OpenStudy (anonymous):

okay, so i have a math packet to do and one of the sections is to graph each system of equations and state its solution. 1.) x+2y=-7 2x-3y=o i dont know how to do this.

OpenStudy (anonymous):

For the top equation, you may first let x=0 and solve for y: x=0 ==> 2y=-7 ==> y=-7/2 ==> (0,-7/2) Then repeat for y=0: y=0 ==> x=-7 ==> (-7,0) The you may use these two pints to graph the line. (cont)

OpenStudy (anonymous):

For the bottom equation you might want to solve for y to get into the slope-intercept form since the x-int and y-int are the same: y=2x/3 Then, you can pick any convenient value of x. A multiple of 3 is good here: x=6 ==> y=2(6)/3=4 ==> (6,4) The other point to use on this line is (0,0). Now you can graph the second line. (cont)

OpenStudy (anonymous):

The intersection of the two lines is the solution. Beware it takes some precision to get the answer this way. Use graph paper if you have it. I will attach the graph if you give me a minute. (cont)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The solution is (-3,-2) and the graph is attached. let me know if you have any questions.

OpenStudy (anonymous):

i just dont understand this AT ALL. i have such a hard time.

OpenStudy (anonymous):

You know that line have the form ax+y=c where a,b,c are numbers? When you have two of them, you have a "system" the solution is where they intersect.

OpenStudy (anonymous):

alright.. i can understand that.

OpenStudy (anonymous):

And you know that it takes two points to "fix" or locate a line in the plane. So what I did was find two points for a each line, and graph them using a software that I have so that I could show you, but it is easily done by hand. Then I looked for the point of intersection which turned out to be (-3,-2).